# VideoGLUE: Video General Understanding Evaluation of Foundation Models

### v1

@@ Abstract
We evaluate existing foundation models’ video understanding capabilities using a carefully designed experiment protocol consisting of three hallmark tasks (action recognition, temporal localization, and spatiotemporal localization), eight datasets well received by the community, and four adaptation methods tailoring a foundation model (FM) for a downstream task. Moreover, we propose a scalar VideoGLUE score (VGS) to measure an FM’s efficacy and efficiency when adapting to general video understanding tasks. Our main findings are as follows. First, task-specialized models significantly outperform the six FMs studied in this work, in sharp contrast to what FMs have achieved in natural language and image understanding. Second, video-native FMs, whose pretraining data contains the video modality, are generally better than image-native FMs in classifying motion-rich videos, localizing actions in time, and understanding a video of more than one action. Third, the video-native FMs can perform well on video tasks under light adaptations to downstream tasks (e.g., freezing the FM backbones), while image-native FMs win in full end-to-end finetuning. The first two observations reveal the need and tremendous opportunities to conduct research on video-focused FMs, and the last confirms that both tasks and adaptation methods matter when it comes to the evaluation of FMs.
@@ 1 Introduction
Foundation models (FMs) are a term coined by Bommasani et al. [7], referring to “any model that is trained on broad data that can be adapted (e.g., finetuned) to a wide range of downstream tasks.” Some representative FMs include but are not limited to BERT [13], GPT-3 [8], CLIP [41], and ALIGN [26]. This work primarily investigates the video understanding capabilies of six visual and multimodal FMs: CoCa [59], CLIP [41], FLAVA [47], VideoMAE [48], VATT [1], and InternVideo [55]. These models are selected because they are amendable for the video understanding of our interest and make their checkpoints accessible to us.
It is nontrivial to evaluate FMs. In contrast to “specialist” models developed for a particular task, FMs are considered as “generalists” that learn shareable meta-knowledge across tasks so that one can quickly adapt them to achieve superior performance on various downstream tasks. Hence, both the tasks and adaptation methods matter when it comes to evaluation. However, the community has not reached a consensus on these two aspects. FM developers select their own different sets of downstream tasks — interestingly, often covering no video or only appearance-rich video classification tasks [9, 30]. Moreover, they rely on distinct adaptation methods, making apples-to-apples comparisons challenging and causing mismatches with the FMs’ actual use cases.
Figure 1: FMs vs. state-of-the-art task-specialized models on video understanding. Unlike natural language and image understanding, video tasks are where FMs generally fall behind “specialists”.
@@
Why do we specifically focus on videos? The main motivation is to promote video understanding in the evaluation of FMs. More concretely, we test the following conjectures through this work. First, FMs’ high performance on existing evaluation suites does not necessarily indicate their potential in video since these suites either lack video-specific tasks or selectively choose video tasks whose appearance feature is more important than motion — InternVideo [55] is an exception as discussed in the next paragraph. Second, existing FMs cannot heed motion in video, given that they learn primarily from static images (and corresponding text descriptions) or short video clips containing limited motion. Third, popular adaptation methods (e.g., finetuning all weights) cannot supplement FMs with all the cues needed to recognize motion-rich actions and localize entities temporally and/or spatiotemporally.
While our work is not the first to emphasize the evaluation of FMs, it is unique on multiple fronts. Unlike Elevater [32]’s target of evaluating language-augmented FMs, we consider all FMs adaptable to video understanding. Unlike Perception Test [4]’s coverage of a broad spectrum of perception tasks, we focus on video, allowing us to cover various aspects of this vertical domain. Interestingly, many of our datasets also appear in InternVideo [55], a video-oriented FM. However, we promote model adaptation methods as an inherent part of the evaluation protocol — a consistent set of diverse adaptation methods is necessary to provide FMs ample opportunities to expose their video understanding capabilities. Moreover, unlike InternVideo’s focus on their single FM, we evaluate FMs developed by different research groups in an uniform experiment protocol — the first of its kind for visual and multimodal FMs, to the best of our knowledge.
Our main findings are as follows. First, task-specialized models still significantly outperform the six FMs studied in this work (see Figure 1), in sharp contrast to what FMs have achieved in natural language and image understanding. Hence, there is a need and tremendous opportunities to research video-focused FMs. Second, video-native FMs, whose pretraining data contains the video modality, are generally better than image-native FMs in classifying motion-rich videos, localizing actions in time, and understanding a video of more than one action. Third, the video-native FMs can perform well on video tasks under light adaptations to downstream tasks (e.g., freezing the FM backbones), while image-native FMs win in full end-to-end finetuning. This observation confirms that both tasks and adaptation methods matter when it comes to the evaluation of FMs.
@@ 2 Related work
One common type of FMs are Large Language Models (LLMs) trained to acquire generic, transferable, and diverse representations that can enable sample-efficient learning and knowledge transfer across a broad range of downstream tasks. FMs are often trained with simple self-supervised learning objectives such as predicting the next token in a sentence (e.g., GPT-3 [8], PaLM [12]), or denoising the masked tokens (e.g., BERT [13], UNILM [14], and BEiT [5]). An intriguing characteristic of FMs is their ability to gradually acquire new capabilities as the model grows and the training data size increases, despite being trained on simple learning objectives [56]. For example, PaLM [12, 3], a massive LM with 540 billion parameters has started to show new capabilities in tasks such as explaining jokes, solving math, and performing common-sense reasoning when scaled to over 100B parameters.
@@
In addition to self-supervised transformers, FMs in computer vision also encompass transformers specifically trained to align image-text paired data. These FMs use learning objectives include contrastive learning (e.g., CLIP [41]), denoising masked tokens (e.g., BEiT-3 [53]), predicting the next token in a single modality (e.g., DALL-E [43]) or in the interleaved image-text sequence (e.g., Flamingo, KOSMOS-1 [24]). Recent FMs are also trained on a mixture of these objectives (e.g., CoCa [59], FLAVA [47], MAE [22]). For example, MAE combines autoencoder reconstruction objective jointly with the denoising objective [22] that was extended to video [18, 48]. In our study, we choose six representative FMs (i.e., CoCa [59], CLIP [41], FLAVA [47], VideoMAE [48], VATT [1], and InternVideo [55]) due to their amendability on video understanding and accessibility of checkpoints.
Evaluation of FMs.
@@
As the mission of FMs is to enable sample-efficient knowledge transfer, the design of downstream tasks is critical to evaluate the capabilities and limitations of these models. The evaluation of FMs is pioneered by the NLP researchers. For example, GLUE [50] and SuperGLUE [49] introduced a suite of tools for evaluating language understanding tasks. The authors utilized established public benchmarks and provided tools for evaluating, probing, and benchmarking pretrained FMs, allowing for a comparison to human baselines. ELEVATER [32] introduced this concept to vision FMs along with a toolkit for evaluating vision-language tasks, including knowledge augmentation, hyperparameter tuning, and three adaptation techniques. In parallel, there have been attempts to establish a diagnostic benchmark for perceptual understanding of the world. For instance, Perception Test [4] crowd-sourced 11K videos in which about 100 users performed scripted activities. This benchmark [4] comprises videos filmed by only about 100 participants, which may not provide the same level of domain coverage and diversity as the other FM evaluation works mentioned earlier.
Evaluation of video FMs.
@@
While some vision-language FMs have incorporated video tasks, their evaluation typically follows that of static images and neglects the unique aspects of video spatial-temporal modeling and reasoning. To our knowledge, no previous work has been solely dedicated to evaluating Video FMs. The closest work to ours are InternVideo [55] and VideoMAE [48], which introduce new FMs and show their superiority over several dozen video datasets. There are two key differences to the prior works. First, our evaluation is video-centric using the tasks that require motion understanding or long-term temporal reasoning. Second, instead of promoting new video FMs, our work proposes no new models and is solely dedicated to evaluating current and future Video FMs in an impartial reproducible experimental setup. Concretely, our goal is to provide tools for probing and benchmarking FMs on motion tasks in various setting include using the parameter-efficient adapter.
[…]
Foundation Model Modality Pretraining Data Pretraining Objective
CoCa [59] Image + Text JFT3B [60] + ALIGN [26] Contrastive + Captioning
CLIP [41] Image + Text WebImageText [41] Contrastive
FLAVA [47] Image + Text PMD [47] Contrastive + MIM + MLM
VideoMAE [18] Video K400 [28] MVM
InternVideo [55] Video UnlabeledHybrid [55] MVM + Contrastive
VATT [1] Video + Audio + Text HT100M [37] Contrastive
3 FMs for video understanding
In this paper, we are interested in examining which FMs are good at solving video tasks, what makes them better than others in the video domain, and how to best adapt them to video understanding. Table 1 shows the six FMs we gained access to via public repositories or personal communications.
@@ 3 FMs for video understanding
4 Tasks and adaptation methods both matter to the evaluation of FMs
@@
This section describes our video general understanding evaluation (VideoGLUE) benchmark: video-focused downstream tasks and methods to adapt an FM to the tasks. The former concretizes the video understanding capabilities we want to evaluate from an FM, while the latter provides various paths for an FM to showcase the corresponding capabilities.
@@
Table 2: Summary of statistics, video properties, and data sources of each dataset. Tasks involved are spatiotemporal action localization (STAL), temporal action localization (TAL), and video classification (VC). Column "Num. videos" contains video examples in train/evaluation splits, respectively.
@@
Task Dataset Num. videos Avg. length Data source Note
STAL AVA v2.2 [21] $210,634$ / $57,371$ 15 mins Movie spatiotemporal, instance
AVA-Kinetics [31] $354,201$ / $91,919$ 10 seconds Web spatiotemporal, instance
TAL ActivityNet v1.3 [16] $10,002$ / $4,926$ 5-10 mins Web temporal
VC Kinetics400 [28] $235,693$ / $19,165$ 10 seconds Web holistic, appearance
Moments-in-Time [38] $791,246$ / $33,898$ 3 seconds Web holistic, appearance
Sth-sth v2 [20] $168,913$ / $24,777$ 2-6 seconds Crowd-source holistic, motion
Diving48 [34] $15,027$ / $1,970$ 5 seconds Web holistic, motion
Charades [46] $110,905$ / $4,985$ 30 seconds Crowd-source multi-label, long-clip
@@ 4.1 Video understanding tasks
Like objects’ role in image understanding, actions are the core of video understanding, leading us to select tasks and datasets that recognize and localize actions in time and space. Table 2 provides a quick summary. Next, we explain the rationale behind the particular choices of datasets and postpone the datasets’ details to the supplementary materials.
@@ 4.1.1 Recognizing actions
General actions. We first include the action recognition datasets of Kinetics400 (K400) [28], Moments-in-Time (MiT) [38], and Charades [46], considering their popularity that they are being complementary to each other. Regarding data sources, K400 videos are from Youtube, MiT draws videos from different Web venues, while Charades contains scripted videos. Regarding action labels, the datasets differ in granularities and real-life scenarios, a verb defines an action in MiT, K400 groups actions by verb-subject pairs, and Charades actions are about indoor activities. Regarding the average length, K400 and MiT videos are between 3 and 10 seconds, each with one action label, while Charades videos are about 30 seconds, each with multiple actions.
Fine-grained motion-focused actions. We also include Something-something-v2 (SSv2) [20] and Diving48 (D48) [34] as another two action recognition datasets, whose actions are fine-grained and motion-focused. SSv2 contains 174 human hand gestures as action labels, such as putting something into something, turning something upside down, and covering something with something. D48 is all about competitive diving. Notably, the foreground objects’ motion is a more significant discriminative cue than their appearance.

### v3

@@ Abstract
We evaluate the video understanding capabilities of existing foundation models (FMs) using a carefully designed experiment protocol consisting of three hallmark tasks (action recognition, temporal localization, and spatiotemporal localization), eight datasets well received by the community, and four adaptation methods tailoring an FM for downstream tasks. Furthermore, we jointly profile FMs’ efficacy and efficiency when adapting to general video understanding tasks using cost measurements during both training and inference. Our main findings are as follows. First, task-specialized models significantly outperform the seven FMs studied in this work, in sharp contrast to what FMs have achieved in natural language and image understanding. Second, video-native FMs, whose pretraining data mainly contains the video modality, are generally better than image-native FMs in classifying motion-rich videos, localizing actions in time, and understanding a video of more than one action. Third, the video-native FMs can perform well on video tasks under light adaptations to downstream tasks (e.g., freezing the FM backbones), while image-native FMs win in full end-to-end finetuning. The first two observations reveal the need and tremendous opportunities to conduct research on video-focused FMs, and the last confirms that both tasks and adaptation methods matter when it comes to the evaluation of FMs. Our code is released under https://github.com/tensorflow/models/tree/master/official/projects/videoglue.
@@ 1 Introduction
Foundation model (FM) is a term coined by Bommasani et al. (2021), referring to “any model that is trained on broad data that can be adapted (e.g., finetuned) to a wide range of downstream tasks.” Some representative FMs include but are not limited to BERT (Devlin et al., 2018), GPT-3 (Brown et al., 2020), CLIP (Radford et al., 2021), and ALIGN (Jia et al., 2021). This work primarily investigates the video understanding capabilies of seven visual and multimodal FMs: CLIP (Radford et al., 2021), FLAVA (Singh et al., 2022), CoCa (Yu et al., 2022), DINOv2 (Oquab et al., 2023), VATT (Akbari et al., 2021), VideoMAE (Tong et al., 2022), and InternVideo (Wang et al., 2022b). We select these models because they are amendable for the video understanding and make their checkpoints accessible to us.
It is nontrivial to evaluate FMs. In contrast to “specialist” models developed for a particular task, FMs are considered as “generalists” that learn shareable meta-knowledge across tasks so that one can quickly adapt them to achieve superior performance on various downstream tasks. Hence, both the tasks and adaptation methods matter when it comes to the evaluation of FMs. However, the community has not reached a consensus on these two aspects. FM developers select their own different sets of downstream tasks — interestingly, often covering no video or only appearance-rich video classification tasks (Buch et al., 2022; Lei et al., 2023). Moreover, they rely on distinct adaptation methods, making apples-to-apples comparisons challenging and causing mismatches with the FMs’ actual use cases.
(a) CLIP
(b) FLAVA
(c) CoCa
(d) DINOv2
(e) VATT
(f) VideoMAE
(g) InternVideo
Figure 1: Performance of FMs with end-to-end finetuning (red) and frozen backbone (blue), in comparison with state-of-the-art task-specialized models (black) on VideoGLUE benchmarks. VC-A, VC-M, and VC-ML stand for appearance-focused, motion-focused, and multi-label Video Classification tasks, respectively; TAL stands for Temporal Action Localization; STAL stands for Spatiotemporal Action Localization. The highest and lowest performance numbers on each dataset are mapped to $0.9$ and $0.1$, and the other numbers are linearly scaled accordingly on the radar chart. We also use gray shades to represent tasks that are more focused on appearance understanding more than motion. We observe that: (1) FMs generally fall behind task-specialized models; (2) FMs that are trained with video data are generally better than image-native FMs on motion-focused tasks under the frozen backbone setting, and image-native FMs can generally catch up when finetuned end-to-end on the target dataset.
@@
Why do we specifically focus on videos? The main motivation is to promote video understanding in the evaluation of FMs. More concretely, we test the following conjectures through this work. First, FMs’ high performance on existing evaluation suites does not necessarily indicate their potential in video since these suites either lack video-specific tasks or selectively choose video tasks whose appearance feature is more important than motion — InternVideo (Wang et al., 2022b) is an exception as discussed in the next paragraph. Second, many existing FMs are unlikely to be able to handle motion in video well, given that they learn primarily from static images (Radford et al., 2021; Singh et al., 2022; Yu et al., 2022) or short video clips containing limited motion (Feichtenhofer et al., 2022; Wang et al., 2022b). Third, popular adaptation methods (e.g., finetuning all weights) cannot supplement FMs with all the cues needed to recognize motion-rich actions and localize entities temporally and/or spatiotemporally, as elaborated in Sections 4.1 and 4.2.
While our work is not the first to emphasize the evaluation of FMs, it is unique on multiple fronts. Unlike ELEVATER (Li et al., 2022a)’s target of evaluating language-augmented FMs, we consider all FMs adaptable to video understanding which does not necessarily involve language. Unlike Perception Test (Patraucean et al., 2024)’s coverage of a broad spectrum of perception tasks, we focus on video, allowing us to cover various aspects of this vertical domain. Interestingly, many of our datasets also appear in InternVideo (Wang et al., 2022b), a video-oriented FM. However, we promote model adaptation methods as an inherent part of the evaluation protocol — a consistent set of diverse adaptation methods is necessary to provide FMs ample opportunities to expose their video understanding capabilities. Moreover, unlike InternVideo’s focus on their single FM, we evaluate FMs developed by different research groups in an uniform experiment protocol — the first of its kind for visual and multimodal FMs, to the best of our knowledge.
Our main findings are as follows. First, task-specialized models still significantly outperform the seven FMs studied in this work (see Figure 1), in sharp contrast to what FMs have achieved in natural language (OpenAI, 2022; Roberts et al., 2022) and image understanding (Radford et al., 2021; Yu et al., 2022; Chen et al., 2022). Hence, there is a need and tremendous opportunities to research video-focused FMs. Second, video-native FMs, whose pretraining data mainly contains the video modality, are generally better than image-native FMs in classifying motion-rich videos, localizing actions in time, and understanding a video of more than one action. Third, the video-native FMs can perform well on video tasks under light adaptations to downstream tasks (e.g., freezing the FM backbones), while image-native FMs win in full end-to-end finetuning. This observation confirms that both tasks and adaptation methods matter when it comes to the evaluation of FMs.
@@ 2 Related work
One common type of FMs are Large Language Models (LLMs) trained to acquire generic, transferable, and diverse representations that can enable sample-efficient learning and knowledge transfer across a broad range of downstream tasks. FMs are often trained with simple self-supervised learning objectives such as predicting the next token in a sentence (e.g., GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2022)), or denoising the masked tokens (e.g., BERT (Devlin et al., 2018), UniLM (Dong et al., 2019), and BEiT (Bao et al., 2021)). An intriguing characteristic of FMs is their ability to gradually acquire new capabilities as the model grows and the training data size increases, despite being trained on simple learning objectives (Wei et al., 2022). For example, PaLM (Chowdhery et al., 2022; Anil et al., 2023), a massive LM with $540$ billion parameters, has started to show new capabilities in tasks such as explaining jokes, solving math, and performing common-sense reasoning when scaled to over 100B parameters.
@@
In addition to self-supervised transformers, FMs in computer vision also encompass transformers specifically trained to align image-text paired data. These FMs use learning objectives include contrastive learning (e.g., CLIP (Radford et al., 2021)), denoising masked tokens (e.g., BEiT-3 (Wang et al., 2022a)), predicting the next token in a single modality (e.g., DALL-E (Ramesh et al., 2021)) or in the interleaved image-text sequence (e.g., Flamingo (Alayrac et al., 2022), Kosmos-1 (Huang et al., 2023)). Recent FMs are also trained on a mixture of these objectives (e.g., MAE (He et al., 2022), FLAVA (Singh et al., 2022), and CoCa (Yu et al., 2022)). Our criteria of choosing foundation models to study are primarily based on the definition of FMs, and their amendability on video understanding and accessibility of checkpoints. We leave some models (e.g., detection, segmentation models) out of the scope of this work, because of their current lack of generalization on video understanding tasks. Finally we choose seven representative FMs, i.e., CLIP (Radford et al., 2021), FLAVA (Singh et al., 2022), CoCa (Yu et al., 2022), DINOv2 (Oquab et al., 2023), VATT (Akbari et al., 2021), VideoMAE (Tong et al., 2022), and InternVideo (Wang et al., 2022b).
Evaluation of foundation models.
@@
As the mission of FMs is to enable sample-efficient knowledge transfer, the design of downstream tasks is critical to evaluate the capabilities and limitations of these models. The evaluation of FMs is pioneered by the NLP researchers. For example, GLUE (Wang et al., 2018a) and SuperGLUE (Wang et al., 2019) introduced a suite of tools for evaluating language understanding tasks. The authors utilized established public benchmarks and provided tools for evaluating, probing, and benchmarking pretrained FMs, allowing for a comparison to human baselines. ELEVATER (Li et al., 2022a) introduced this concept to vision FMs along with a toolkit for evaluating vision-language tasks, including knowledge augmentation, hyperparameter tuning, and three adaptation techniques. In parallel, there have been attempts to establish a diagnostic benchmark for perceptual understanding of the world. For instance, Perception Test (Patraucean et al., 2024) crowd-sourced 11K videos in which about 100 users performed scripted activities. This benchmark comprises videos filmed by only about 100 participants, which may not provide the same level of domain coverage and diversity as the other FM evaluation works mentioned earlier.
Evaluation of video foundation models.
@@
While some vision-language FMs have incorporated video tasks, their evaluation typically follows that of static images and neglects the unique aspects of video spatial-temporal modeling and reasoning. To our knowledge, no previous work has been solely dedicated to evaluating video FMs. The closest work to ours are InternVideo (Wang et al., 2022b) and VideoMAE (Tong et al., 2022), which introduce new FMs and show their superiority over several video datasets. This work has two key differences to the prior ones. First, our evaluation is video-centric using the tasks that require motion understanding or long-term temporal reasoning. Second, instead of promoting new video FMs, our work proposes no new models and is solely dedicated to evaluating current and future video FMs in an impartial reproducible experimental setup. Concretely, our goal is to provide tools for probing and benchmarking FMs on motion tasks in various settings.
Table 1: Foundation models studied in this work (MxM stands for Masked Image/Language/Video Modeling).
Foundation Model Pretraining Modality Pretraining Data Pretraining Objective
CoCa (Yu et al., 2022) Image + Text JFT3B + ALIGN Contrastive + Captioning
CLIP (Radford et al., 2021) Image + Text WebImageText Contrastive
FLAVA (Singh et al., 2022) Image + Text PMD Contrastive + MIM + MLM
DINOv2 (Oquab et al., 2023) Image LVD-142M MIM + DINO
VideoMAE (Tong et al., 2022) Video K400 MVM
InternVideo (Wang et al., 2022b) Video UnlabeledHybrid MVM + Contrastive
VATT (Akbari et al., 2021) Video + Audio + Text HT100M Contrastive
@@
3 Tasks and adaptation methods both matter when evaluating foundation models
@@
This section describes our video general understanding evaluation (VideoGLUE) benchmark. We first introduce the visual and multimodal FMs evaluated in this work. Then we discuss the video-focused downstream tasks and methods to adapt an FM to the tasks. The former concretizes the video understanding capabilities we want to evaluate from an FM, while the latter provides various paths for an FM to showcase the corresponding capabilities.
3.1 Foundation models for video understanding
We are interested in examining which FMs are good at solving video tasks, what makes them better than others in the video domain, and how to best adapt them to video understanding. Table 1 shows the seven FMs we gained access to via public repositories or personal communications. Thanks to the powerfulness and scalability of the transformer architecture (Vaswani et al., 2017), most developed FMs converge to adopt the vision transformer architecture. Thus for all evaluated FMs, we intentionally choose the ViT-B (Dosovitskiy et al., 2020) variant to bring fair comparison into our benchmark. We also notice, in previous literature, models may be evaluated with different number of frames and resolutions, resulting in unfair comparison (Yan et al., 2022; Feichtenhofer et al., 2021). In VideoGLUE, we control the number of tokens observed by the ViT to be consistent for different FMs on the same task. For the detailed setup on each dataset, please refer to supplementary materials D.
@@ 3.1 Foundation models for video understanding
Table 2: Summary of statistics, video properties, and data sources of each dataset. Tasks involved are video classification (VC), spatiotemporal action localization (STAL), and temporal action localization (TAL).
@@
Task Dataset # of videos (train/validation) Avg. length Source Notes
VC Kinetics-400 $235,693$ / $19,165$ $10$ secs Web Holistic, appearance
Moments in Time $791,246$ / $33,898$ $3$ secs Web Holistic, appearance
Something-Something v2 $168,913$ / $24,777$ $2\sim 6$ secs Crowdsource Holistic, motion
Diving48 $15,027$ / $1,970$ $5$ secs Web Holistic, motion
Charades $7,811$ / $1,814$ $30$ secs Crowdsource Multi-label, long-clip
TAL ActivityNet v1.3 $10,002$ / $4,926$ $5\sim 10$ mins Web Temporal
STAL AVA v2.2 $210,634$ / $57,371$ $15$ mins Movie Spatiotemporal, instance
AVA-Kinetics $354,201$ / $91,919$ $10$ secs Web Spatiotemporal, instance
@@ 3.2 Video understanding tasks
Like objects’ role in image understanding, actions are the core of video understanding, leading us to select tasks and datasets that recognize and localize actions in time and space. Table 2 provides a quick summary. Next, we explain the rationale behind the particular choices of datasets and postpone the datasets’ details to the supplementary materials B.
@@ 3.2.1 Recognizing actions
General actions. We first include the action recognition datasets of Kinetics-400 (K400) (Kay et al., 2017), Moments in Time (MiT) (Monfort et al., 2019), and Charades (Sigurdsson et al., 2016), considering their popularity and they are complementary to each other. Data domain coverage is an important factor when designing benchmarks for FMs, as nowadays FMs are typically trained on massive data sources. K400 videos are from YouTube, MiT draws videos from different Web venues, while Charades contains scripted indoor videos. Internet often returns entertaining and atypical videos, while Charades is about typical everyday videos (Sigurdsson et al., 2016). Regarding action labels, the datasets differ in granularities and real-life scenarios: a verb defines an action in MiT, K400 groups actions by verb-subject pairs, and Charades actions are about indoor activities. Regarding the average length, K400 and MiT videos are between 3 and 10 seconds, each with one action label, while Charades videos are about 30 seconds, each with multiple actions.
Fine-grained motion-focused actions. We also include Something-something v2 (SSv2) (Goyal et al., 2017) and Diving48 (D48) (Li et al., 2018) as another two action recognition datasets, whose actions are fine-grained and motion-focused. SSv2 contains 174 human hand gestures as action labels, such as putting something into something, turning something upside down, and covering something with something. The videos are mostly focusing on hand-object interactions. D48 videos are all about competitive diving recordings collected from Web sources. Notably, in these datasets the foreground objects’ motion is a more significant discriminative cue than their appearance.